Proceedings of 1993 Oberwolfach Conference Novikov Conjectures, Index Theorems and Rigidity edited by S. Ferry, A. Ranicki and J. Rosenberg LMS Lecture Notes Vols. 226, 227, Cambridge (1995)

Contents, Volume 1

Preface, The Editors . . . . . . . . . . . . . . . . . . . . . vii Introductory Material Programme of the Conference . . . . . . . . . . . . . . . . . . . 1 Participant List from the Conference . . . . . . . . . . . . . . . 3 A history and survey of the Novikov conjecture,

Steven C. Ferry, Andrew Ranicki, and Jonathan Rosenberg …. 7 Annotated Problem List . . . . . . . . . . . . . . . . . . . . 67 Classic Papers On the Steenrod homology theory (first distributed 1961),

John Milnor . . . . . . . . . . . . . . . . . . . . . . . 79 Homotopy type of differentiable manifolds (first distributed 1962),

William Browder . . . . . . . . . . . . . . . . . . . . . 97 K-theory, group C^{∗}-algebras, and higher signatures (Conspectus)

(first distributed 1981), Gennadi G. Kasparov . . . . . . . . 101 Research and Survey Papers A coarse approach to the Novikov conjecture,

Steven C. Ferry and Shmuel Weinberger . . . . . . . . . . . 147 Geometric reflections on the Novikov conjecture, Mikhael Gromov . . 164 Controlled Fredholm representations, Alexander S. Mishchenko … 174 Assembly maps in bordism-type theories, Frank Quinn . . . . . . . 201 On the Novikov conjecture, Andrew Ranicki . . . . . . . . . . . 272 Analytic Novikov for topologists, Jonathan Rosenberg . . . . . . . 338

Contents, Volume 2

Preface, The Editors . . . . . . . . . . . . . . . . . . . . . vii Research and Survey Papers Proper affine isometric actions of amenable groups,

M. E. Bekka, P.-A. Cherix, and Alain Valette . . . . . . . . . . 1 Bounded K-theory and the assembly map in algebraic K-theory,

Gunnar Carlsson . . . . . . . . . . . . . . . . . . . . . . 5 On the rigidity of the index on open manifolds, Jürgen Eichhorn . . 128 Remarks on Steenrod homology, Steven C. Ferry . . . . . . . . . 148 Epsilon surgery theory, Steven C. Ferry and Erik K. Pedersen … 167 On the coarse Baum-Connes conjecture, Nigel Higson and John Roe . 227 Exotic index theory and the Novikov conjecture, Steven Hurder … 255 Bounded and continuous control, Erik K. Pedersen . . . . . . . . 277 On the homotopy invariance of the boundedly controlled signature of a

manifold over an open cone,

Erik K. Pedersen, John Roe, and Shmuel Weinberger . . . . . 285 3-manifolds and PD(3)-groups, C. B. Thomas . . . . . . . . . . 301 Orthogonal complements and endomorphisms of Hilbert modules and

C^{∗}-elliptic complexes, Evgenii V. Troitsky . . . . . . . . . . . 309 Assembly, Michael Weiss and Bruce Williams . . . . . . . . . . . 332 Pro-excisive functors, Michael Weiss and Bruce Williams . . . . . . 353